OpenStudy (anonymous):

Put the following numbers in order from smallest to largest. 0.085 1/3 2/5 60% 3/4 78% 1.05

OpenStudy (anonymous):

@Ashleyisakitty @beccaboo333 @Compassionate @djcool31 @e.mccormick @fathomboy @Grazes @HelpBlahBlahBlah @iPwnBunnies @jdoe0001 @kelci2014 @Luigi0210 @mathmale @nikki__kitty @oneinmillon @phi @kelci2014 @Rea201 @Squirrels @thomaster @utkFRESH @violetcateyes @Whitemonsterbunny17 @utkFRESH @Xmoses1 @zubhanwc3 sorry for the spam but can anyone help me?

OpenStudy (anonymous):

use a calculator for the fractions, change the % to /100 (divide by 100) ex. 3/4=.75 50%=50/100=.5 Then, compare the numbers and change them into the right order

OpenStudy (anonymous):

0.085, 1/3, 2/5, 60%, 3/4, 78%, 1.05 They appear to already be in order from smallest to largest.

OpenStudy (anonymous):

@Grazes is she right?

OpenStudy (pinksapphire):

She is right.

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

But you should do as I prescribed just to check if you can't mentally check it quickly.

OpenStudy (e.mccormick):

A common denominator form works for these, and as Grazes points out, a calculator and decimal form is good for most times you see these questions. It really is a common denominator at that point.